Find three ordered pairs that satisfy the equation.
y = 2x − 4
(x, y) = 0, _____
(x, y) = 1, _____
(x, y) = 2, _____




Answer by richwmiller(17219)

plug in those numbers for x and calculate y
y = 2*0 − 4
(x, y) = 0, _____
y = 2*1 − 4
(x, y) = 1, _____
y = 2*2 − 4
(x, y) = 2, _____
Certainly, you can do the rest.